A.The laser system tested for inertial confinement can produce a 109.6 kJ pulse only 1.1 ns in duration. How many photons (in terms of 1023 photons) are in the pulse if their wavelength is 0.97 μm?

B.Calculate the energy in MeV released in a α decay of 249Cf. Assume that the masses of 249Cf and 245Cm are 249.077455 and 245.060938 u, respectively.

C.Calculate the energy in MeV released in a β- decay of 90Sr. Assume that the masses of 90Sr and 90Y are 89.909088 and 89.90655 u, respectively.

D.The detail observable using a probe is limited by its wavelength. Calculate the energy in nanojoules of a γ-ray photon that has a wavelength of 2.84×10−16 m.
Answer #1

ONE QUESTION AT A TIME, PLEASE

A)

Energy, E = hc / \lambda = 6.63e-34 * 3e8 / 0.97e-6 = 2.0505e-19 J

so,

Number of photons = 109.6e3 / 2.114e-19

Number of photons = 5.345e23 photons

____________________________________

B)

we need to find mass defect

\Deltam = m (249 Cf) - m ( 245 Cm) - m (4He)

\Deltam = 249.077455 - 245.060938 - 4.002603

\Deltam = 0.013914 u

so,

E = \Delta m * 931.5

E = 12.9 MeV